Solve the system by the sustitution method: 3x - 2y = 0
5x + 10y =4

Answer by stanbon(75887)

Solve the system by the sustitution method:
3x - 2y = 0
5x + 10y =4
------------
Solve the 1st equation for "y":
y = (3/2)x
-----
Substitute for "y" in the 2nd equation and solve for "x":
5x + 10(3/2)x = 4
5x + 15x = 4
20x = 4
x = 4/20
x = 1/5
===========
Since y = (3/2)x, y = (3/2)(1/5) = 3/10
========================================
Cheers,
Stan H.
===========